ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57P99, 57N35, 57N40 Mio, Washington, Florida State University, USA On the geometric topology of homology manifolds. The discovery of nonresolvable ANR homology manifolds in dimensions ≥ 6 has raised a multitude of questions about the geometric topology of these spaces. We will discuss recent advances that include transversality, embedding, and control improvement theorems.